the average person in the country discards 5 pounds of trash each day. The population of the United States is about 290,000,000 people. How much trash does the population of the country discard in a year? Answer in scientific notation.
1
Expert's answer
2012-12-11T09:28:44-0500
If the average person in the country discards 5 pounds of trash each day, then the population of about 290,000,000 people will discard about

P = 5 · 29·10^7 · 365 = 52925·10^7

pounds of trash in a year.